# EmoSteer-TTS: Fine-Grained and Training-Free Emotion-Controllable Text-to-Speech via Activation Steering

### v2

@@ Conclusion
For all three TTS models used in our study, i.e., F5-TTS (Chen et al. 2025), E2-TTS (Eskimez et al. 2024), and CosyVoice2 (Du et al. 2024), steering operations are implemented as hook functions. These hooks are registered either before or after the forward pass of the first residual stream in each DiT block.
@@ A.4 Multiple Emotion Steering
Appendix B: Dataset Construction
To ensure the effectiveness of the steering vectors, we curate an emotional speech dataset by collecting and filtering audio samples with clearly distinguishable emotional tones from multiple existing corpora, including MSP-Podcast (Lotfian and Busso 2017), IEMOCAP (Busso et al. 2008), RAVDESS (Livingstone and Russo 2018), CREMA-D (Cao et al. 2014), TESS (Pichora-Fuller and Dupuis 2020), SAVEE (Jackson and Haq 2014), ASVP-ESD (Landry et al. 2020), CASIA (CASIA 2023), M3ED (Zhao et al. 2022), ESD (Zhou et al. 2022), and Emo-Emilia (Zhao et al. 2025). The quality filtering process involves the following steps:
@@
    We use librosa (McFee et al. 2015) to remove utterances that are either too short (<<2s) or too long (>>20s).
@@
The resulting dataset covers a broad range of speakers, emotions, and speaking styles, providing a robust foundation for learning and evaluating fine-grained emotion steering in text-to-speech synthesis.
Appendix C: Configurations
@@ C.1 Model Configurations
We steer three pretrained conditional flow matching (CFM)-based TTS models, i.e., F5-TTS, E2-TTS, and CosyVoice2, in our main experiments. As illustrated in Fig.6, we apply steering to the first residual stream at each layer of these models. The detailed configurations of the models are provided in Table 2.
Figure 6: A DiT block in a CFM-based TTS model.
@@ C.2 Hardware and Software Configurations
All experiments were conducted on a server equipped with 8× NVIDIA RTX 6000 Ada GPUs (48GB each) and 2× Intel(R) Xeon(R) Platinum 8375C CPUs (2.9GHz, 32 cores each), with a total of 256GB of RAM. The operating system was Ubuntu 20.04.6 LTS. All code was executed in Conda environments. The relevant software libraries and frameworks for each model (F5-TTS, E2-TTS, CosyVoice2) are described in their GitHub repositories [^1] https://github.com/SWivid/F5-TTS [^2] https://github.com/lucidrains/e2-tts-pytorch [^3] https://github.com/FunAudioLLM/CosyVoice.
Appendix D: Objective Evaluation Metrics
@@ D.3 Emotion Erasure Mean Opinion Score
  • 5Emotion fully removed: The target emotion is completely erased; the speech sounds emotionally neutral and natural, with no detectable emotional cues.
  • 4Emotion mostly removed: Only faint traces of the original emotion remain; the speech is close to neutral.
  • 3Emotion partially removed: The emotional intensity is reduced, but the target emotion is still clearly noticeable.
  • 2Emotion barely removed: The emotional expression remains strong; only minimal reduction is observed.
  • 1Emotion not removed: The original emotional tone persists fully or is even unintentionally enhanced.
@@
Appendix E: Visualization of F0 Contours
E.1 Emotion Interpolation
@@ E.2 Emotion Erasure
In both emotion interpolation and erasure, F0 contours capture only a partial aspect of human emotional perception, as pitch alone cannot fully convey complex emotional nuances. Therefore, we encourage readers to listen to the audio samples available on our demo page for a more complete experience.

### v3

@@ Appendix A Appendix A: Code Snippets for Fine-Grained Emotion Control
For all three TTS models used in our study, i.e., F5-TTS (Chen et al., 2025), E2-TTS (Eskimez et al., 2024), and CosyVoice2 (Du et al., 2024), steering operations are implemented as hook functions. These hooks are registered either before or after the forward pass of the first residual stream in each DiT block. Code will be released upon acceptance.
@@ A.4 Multiple Emotion Steering
Appendix B Appendix B: Dataset Construction
To ensure the effectiveness of the steering vectors, we curate an emotional speech dataset by collecting and filtering audio samples with clearly distinguishable emotional tones from multiple existing corpora, including MSP-Podcast (Lotfian & Busso, 2017), IEMOCAP (Busso et al., 2008), RAVDESS (Livingstone & Russo, 2018), CREMA-D (Cao et al., 2014), TESS (Pichora-Fuller & Dupuis, 2020), SAVEE (Jackson & Haq, 2014), ASVP-ESD (Landry et al., 2020), CASIA (CASIA, 2023), M3ED (Zhao et al., 2022), ESD (Zhou et al., 2022), and Emo-Emilia (Zhao et al., 2025). The quality filtering process involves the following steps:
@@ Appendix B Appendix B: Dataset Construction
    We use librosa (McFee et al., 2015) to remove utterances that are either too short (<<2s) or too long (>>20s).
@@
The resulting dataset covers a broad range of speakers, emotions, and speaking styles, providing a robust foundation for learning and evaluating fine-grained emotion steering in text-to-speech synthesis. Data and code will be released upon acceptance.
Appendix C Appendix C: Configurations
@@ C.1 Model Configurations
We steer three pretrained conditional flow matching (CFM)-based TTS models, i.e., F5-TTS, E2-TTS, and CosyVoice2, in our main experiments. As illustrated in Fig.8, we apply steering to the first residual stream at each layer of these models. The detailed configurations of the models are provided in Table 2. emotion2vec and SenseVoice checkpoints are downloaded from their official repos [^1] https://huggingface.co/emotion2vec/emotion2vec_plus_large [^2] https://huggingface.co/FunAudioLLM/SenseVoiceSmall.
@@ C.2 Hardware and Software Configurations
All experiments were conducted on a server equipped with 8× NVIDIA RTX 6000 Ada GPUs (48GB each) and 2× Intel(R) Xeon(R) Platinum 8375C CPUs (2.9GHz, 32 cores each), with a total of 256GB of RAM. The operating system is Ubuntu 20.04.6 LTS. All code was executed in Conda environments. The relevant software libraries and frameworks for each model (F5-TTS, E2-TTS, CosyVoice2) are described in their GitHub repositories [^3] https://github.com/SWivid/F5-TTS [^4] https://github.com/lucidrains/e2-tts-pytorch [^5] https://github.com/FunAudioLLM/CosyVoice.
Appendix D Appendix D: Objective Evaluation Metrics
@@ D.3 Emotion Erasure Mean Opinion Score
  • 5 — Emotion fully removed: The target emotion is completely erased; the speech sounds emotionally neutral and natural, with no detectable emotional cues.
  • 4 — Emotion mostly removed: Only faint traces of the original emotion remain; the speech is close to neutral.
  • 3 — Emotion partially removed: The emotional intensity is reduced, but the target emotion is still clearly noticeable.
  • 2 — Emotion barely removed: The emotional expression remains strong; only minimal reduction is observed.
  • 1 — Emotion not removed: The original emotional tone persists fully or is even unintentionally enhanced.
@@
Table 3: Unguaranteed reproduced results of open-source baselines.
Method Conversion ($\alpha=2.0$) Interpolation Erasure ($\beta=2.5$)
WER($\downarrow$) S-SIM($\uparrow$) E-SIM($\uparrow$) N-MOS($\uparrow$) EI-MOS($\uparrow$) E-SIM($\uparrow$) $E$-MOS($\uparrow$)
emotion2vec / SenseVoice emotion2vec / SenseVoice
In-distribution evaluation on MSP-Podcast (25% en) and ESD (25% en, 50% zh)
Label-based*
EmoSphere++ 37.29 0.21 0.14 / 0.11avg=0.125 2.14±0.91 2.41±0.83 - -
EmoDubber 65.93 0.16 0.08 / 0.05avg=0.065 1.07±0.94 1.13±1.02 - -
Description
-based*
EmoVoice 5.31 0.48 0.22 / 0.19avg=0.205 3.26±1.22 - - -
CosyVoice2 2.71 0.69 0.23 / 0.25avg=0.240 3.66±1.17 - - -
Unsteered F5-TTS 2.14 0.66 0.07 / 0.04avg=0.055 3.79±0.89 - 0.03 / 0.05avg=0.040 1.21±1.17
E2-TTS 2.71 0.64 0.05 / 0.08avg=0.065 3.51±0.94 - 0.06 / 0.02avg=0.040 1.35±1.05
EmoSteer-TTS#
+ F5-TTS 2.79 0.64 0.29 / 0.26avg=0.275 3.29±1.05 4.00±0.89 0.27 / 0.25avg=0.260 4.02±0.85
+ E2-TTS 3.28 0.59 0.28 / 0.28avg=0.280 3.31±0.97 3.38±1.09 0.24 / 0.26avg=0.250 3.63±1.17
+ CosyVoice2 2.83 0.65 0.26 / 0.29avg=0.275 3.65±1.08 3.56±1.15 0.26 / 0.25avg=0.255 3.94±0.97
Cross-datasets (OOD) evaluation on EMNS (25% en) and SeedTT test sets (25% en, 50% zh)
EmoSteer-TTS#
+ F5-TTS 2.65 0.65 0.25 / 0.27avg=0.260 3.58±1.04 3.46±1.08 0.25 / 0.22avg=0.235 3.92±0.99
+ E2-TTS 3.41 0.55 0.26 / 0.25avg=0.255 3.44±1.07 3.50±0.97 0.24 / 0.27avg=0.255 3.57±1.03
+ CosyVoice2 2.86 0.66 0.28 / 0.25avg=0.265 3.49±1.01 3.48±1.27 0.23 / 0.21avg=0.220 3.98±0.94
*: Training-based, #: Training-free, -: Unsupported operation.
The top three results are indicated in boldface. Unsteered backbones are shown in gray for reference.
@@
Appendix E Appendix E: Reproduced Baseline Results
This section recomputes baselines under a controlled protocol, using the same text prompts, reference speeches, and evaluation scripts as in the in-distribution evaluation on MSP-Podcast and ESD. We report results only for open-source methods, as the reproduced quality cannot be guaranteed. Therefore, the results in Table 3 are provided for reference only.
Appendix F Appendix F: Visualization of F0 Contours
F.1 Emotion Interpolation
@@ F.2 Emotion Erasure
In both emotion interpolation and erasure, F0 contours capture only a partial aspect of human emotional perception, as pitch alone cannot fully convey complex emotional nuances. Therefore, we encourage readers to listen to the audio samples available on our demo page.
Appendix G Appendix G: The Use of LLMs
Some portions of this paper were paraphrased or refined with the assistance of ChatGPT and Gemini. No content was directly generated by LLMs.